#3. A 250-ohm resistor, a 0.450 H inductor, and a 6.45 uF capacitor are connected in series
across an emf with a 36.0 volt amplitude and an angular frequency of 260 rad/s.
a. What is the impedance?
b. What is the current amplitude?
c. What is the phase angle between the voltage and current? Does the voltage lag or lead?
d. What are the voltage amplitudes across the resistor, inductor and capacitor individually?
e. What is the power?
f. What is the resonant frequency of the circuit?
g. What is the impedance at the resonant frequency?
h. What is the power at resonant frequency?
Answer #1

R=2500, L=0 450H c=6.45 HF Vo= 36 v W =260 randl a) Impedance, za JR²+ (%-X.)² XL = LW= Xe=1/W 0.450*260 = 117.12 1 - 5962 6.